OpenStudy (anonymous):

2. Given that the points A, B and C have the coordinates A (0, 6), B (5, 11) and C (2, 3). Find the gradients of AB, AC and BC. Are the three points collinear?

OpenStudy (amistre64):

co-linear? not even close

OpenStudy (amistre64):

the gradients/slopes are point-point and put y/x

OpenStudy (amistre64):

0, 6 -5,-11 ------ -5, -5; slope = -5/-5 = 1

OpenStudy (amistre64):

2,3 -0,-6 ------ 2,-3 ; slope = y/x = -3/2

OpenStudy (amistre64):

5,11 -2,-3 ------ 3,8 ; slope = y/x = 8/3

OpenStudy (amistre64):

since they have different slopes; they are not linear at all

OpenStudy (anonymous):

oh okay thanks so much!